Bunnell Corporation is a manufacturer that uses job-order costing. On January 1, the companys inventory balances were as fol9. Is manufacturing overhead underapplied or overapplied for the year? By how much?2. What is the ending balance in Raw Materials? Raw Materials Beg. Bal. End. Bal.4. What is the total amount of manufacturing overhead applied to production during the year? Manufacturing overhead applied5. What is the total manufacturing cost added to Work in Process during the year? Total manufacturing cost
Answer #1

2. Ending Raw materials = Beginning raw materials + Raw materials purchased - raw materials used in production

=72500+ 628000- 598000= 102500

4. Manufacturing overhead applied to production = predetermined overhead rate * actual direct labor hours worked

=15.50*41,000=635500

5. Total manufacturing cost = Raw materials used + Direct labor + manufacturing overhead applied

= 598,000+ 570,000+ 635,500=1803500

8. Actual manufacturing overhead= Indirect labor + Depreciation, insurance, utilities etc = 150,000+ 470,000= 620,000

9. Manufacturing overhead under applied or over applied

a.Total manufacturing overhead applied =635500

b. Actual manufacturing overhead =620,000

Difference is (a-b)= 15,500

Manufacturing Overhead over applied 15,500

Because applied overhead is more than actual Overhead.